OpenStudy (surana):

A regular octagon has side length 10.9 inches. The perimeter of the octagon is 87.2 in centimeters and the area is 392.4 inches squared. A second octagon has side lengths equal to 21.8 inches. Find the area of the second octagon. Round to the nearest hundredth.

OpenStudy (surana):

I just want to make this clear that question above was asked by someone affiliated with me, and is not a question I'm directly asking. It's just that they're asking me to ask it.

OpenStudy (rational):

you want to learn this and then planning to teach them ?

OpenStudy (surana):

They kind of need help, so I'm intending on learning it and then working it through with them, so basically that.

OpenStudy (rational):

Awesome! start by finding the "scale factor"

OpenStudy (surana):

Scale factor/

OpenStudy (surana):

That's the what one number on the octagon is on the other, right?

OpenStudy (rational):

Exactly! you can obtain "scale factor" by taking ratio of corresponding sides

OpenStudy (michele_laino):

for example, we have: scale factor = 21.8/10.9=...?

OpenStudy (surana):

I think that's a 2/1 ratio, if I'm correct.

OpenStudy (michele_laino):

yes!

OpenStudy (surana):

Okay, so as a 2-1 ratio, given what's going on, I just take 392.4 inches and double it, right?

OpenStudy (michele_laino):

that's right!

OpenStudy (rational):

thats a very good wrong guess, but you're very close!

OpenStudy (surana):

Dah. Let me try again...

OpenStudy (rational):

first square the "scale factor", then multiply it by 392.4

OpenStudy (surana):

Square the scale factor, hang on.

OpenStudy (rational):

you're correct about scale factor = 2

OpenStudy (surana):

475.24/118.81?

OpenStudy (rational):

thats right, but you could have simply squared 2 to get 4

OpenStudy (rational):

scale factor = 2 (scale factor)^2 = 4

OpenStudy (surana):

I thought you meant to choose the actual numbers so that's what I did.

OpenStudy (rational):

Ahh okay i see.. multiply that by 392.4 and you're done

OpenStudy (surana):

So 392.4 x 2 is what I have to do?

OpenStudy (rational):

392.4 x 4 is what you have to work

OpenStudy (surana):

Oh, thank you.

OpenStudy (michele_laino):

Sorry I think about perimeters not area! @Surana

OpenStudy (rational):

basically all regular octagons are similar, and the area scales as "square of scale factor" so when you're given area of some polygon and asked to find the area of another similar polygon, you may work it by simply multiplying the area by (scale factor)^2

OpenStudy (surana):

Okay, thank you. I'll be back in a few moments to see if the person got it.

OpenStudy (surana):

And they did, joy. Thanks everyone! Medals all around.

OpenStudy (rational):

Sweet :)

OpenStudy (michele_laino):

thanks! and sorry again, for my error! @Surana

OpenStudy (surana):

It's alright.